The area of an equilateral triangle is 49 times the square root of 3. Compute the perimeter of this triangle.
Answer by lwsshak3(11628)

The area of an equilateral triangle is 49 times the square root of 3. Compute the perimeter of this triangle.
***
let x=side of equalateral triangle
let h=height
3x=perimeter
..
by pythagorean theorem:
h=√(x^2-(x/2)^2)=√(x^2-x^2/4)=√(3x^2/4)=√3x/2
Area=1/2*base*height=1/2*x*√3x/2=49√3
√3x^2/4=49√3
x^2=4*49
x=2*7=14
3x=42
perimeter of this triangle=42 units